# Attractor learning for spatiotemporally chaotic dynamical systems using echo state networks with transfer learning

### v1

@@ Abstract
In this paper, we explore the predictive capabilities of echo state networks (ESNs) for the generalized Kuramoto-Sivashinsky (gKS) equation, an archetypal nonlinear PDE that exhibits spatiotemporal chaos. We introduce a novel methodology that integrates ESNs with transfer learning, aiming to enhance predictive performance across various parameter regimes of the gKS model. Our research focuses on predicting changes in long-term statistical patterns of the gKS model that result from varying the dispersion relation or the length of the spatial domain. We use transfer learning to adapt ESNs to different parameter settings and successfully capture changes in the underlying chaotic attractor.
Keywords: echo state networks, reservoir computing, generalized Kuramoto-Sivashinsky equation, transfer learning
@@ 1 Introduction
In this paper, we explore the predictive capabilities of echo state networks (ESNs) in the context of spatiotemporal chaos. An echo state network is a simple reservoir computing architecture with the echo state property [28, 29, 39, 7, 30, 21, 58]. We introduce a novel methodology that integrates ESNs with transfer learning (TL). Our goal is to demonstrate that this novel methodology significantly enhances predictive performance when predicting changes in long-term statistical behavior that result from varying model parameters.
@@
in the weak-$\ast$ topology. In this setting, we ask two questions. First, if we fix a parameter value $\gamma_{1}\in\Gamma$ and train an echo state network using trajectories from $\Phi_{\gamma_{1}}$, can the echo state network then predict statistical properties of $(f\circ\Phi_{\gamma_{1}}^{t})_{t\geqslant 0}$, where $f:M\to\mathbb{R}^{n}$ is a given observable? Second, assuming an affirmative answer to the first question, how might we predict the statistical properties of $(f\circ\Phi_{\gamma_{2}}^{t})_{t\geqslant 0}$ with respect to a different parameter value $\gamma_{2}\in\Gamma$?
The point of this paper is to answer these two questions. Our approach is as follows. Suppose the echo state network has been trained using trajectories from $\Phi_{\gamma_{1}}$ and can successfully predict statistical properties of $(f\circ\Phi_{\gamma_{1}}^{t})_{t\geqslant 0}$. Since invariant measures of dynamical systems can depend sensitively on parameters, the invariant measure $\mu_{\gamma_{2}}$ associated with a different parameter value $\gamma_{2}$ might differ meaningfully from $\mu_{\gamma_{1}}$. Consequently, we would not expect the echo state network to successfully predict statistical properties of $(f\circ\Phi_{\gamma_{2}}^{t})_{t\geqslant 0}$ when trained using trajectories from $\Phi_{\gamma_{1}}$. We conjecture that the predictive capability of the echo state network can be ‘transferred’ from the $\gamma_{1}$-system to the $\gamma_{2}$-system using transfer learning [27, 11]. That is, the echo state network will successfully predict statistical properties of $(f\circ\Phi_{\gamma_{2}}^{t})_{t\geqslant 0}$ after we update its training using a small new dataset from the new regime.
@@
Transfer learning is a popular concept in many areas of machine learning, including speech recognition and image processing. However, transfer learning for echo state networks has received comparatively little attention in the literature. We conclude that transfer learning can allow echo state networks to maintain predictive power as model parameters change. Such flexibility is important when statistical properties of the dynamics depend meaningfully on model parameters.
@@ 2.1 Background
with the periodic boundary condition $u(x+L,t)=u(x,t)$ for all $x$ and every $t\geqslant 0$. The KS equation is well-behaved as an infinite-dimensional dynamical system. In particular, the KS model admits a finite-dimensional attractor and is therefore in some sense equivalent to a finite-dimensional dynamical system [13, 20, 43]. See [52] for an overview. Moreover, the KS model admits an inertial manifold that contains the global attractor [32, 12, 53]. The KS equation has been studied extensively, both analytically and computationally [35, 26, 50, 41].
The parameter $L$ (the length of the spatial domain) plays the role of a bifurcation parameter for the KS model. This parameter determines the number of linearly unstable Fourier modes, $S_{\mathrm{unst}}$. In addition, the maximum of the coefficient of the linearized equation occurs at $M_{\mathrm{unst}}$. We refer to $M_{\mathrm{unst}}$ as ”the most unstable wavenumber”. These quantities are given by
@@
[…]
@@ 3 Echo state networks and transfer learning
The input matrix $W_{\mathrm{in}}$ is generated component-wise from a uniform distribution on $[-\beta_{1},\beta_{1}]$. The adjacency matrix is generated as $A=\beta_{2}W_{0}/\lambda$, where $W_{0}$ is a sparse matrix (typically turning on less than $10\%$ of the possible connections), $\lambda$ denotes the largest eigenvalue of $W_{0}$, and $\beta_{2}$ is a scaling parameter.
@@
where $\mathbf{X}$ is the training data (trajectories), and $\mathbf{R}$ denotes the reservoir timeseries (with $A$ and $W_{\mathrm{in}}$ fixed). $W_{\mathrm{out}}$ can be expressed explicitly as
@@
Here, $\mathbf{R}^{\prime}$ denotes the transpose of $\mathbf{R}$. Throughout this paper, we use ESN parameters $D=5000$, $\beta_{1}=0.01$, $\beta_{2}=0.1$, and $\mu=0.000005$, unless stated otherwise.
Transfer learning [27, 11] is a machine-learning approach wherein knowledge in a source domain is used to improve predictive performance in a target domain. In our context, the source and target domains are simply two sets of gKS parameters. Our central premise for linking transfer learning to ESNs is the following. Once the ESN is well trained in the source domain, it will retain predictive power in the target domain after receiving a training update using a small transfer-learning dataset from the target domain.
Writing $\mathbf{X}_{TL}$ for the transfer-learning dataset, the update to the output layer is obtained by solving the optimization problem given by
$\delta W=\operatorname*{arg\,min}_{\delta W}\left\|(W_{\mathrm{out}}+\delta W)\mathbf{R}_{TL}-\mathbf{X}_{TL}\right\|_{2}^{2}+\alpha\|\delta W\|_{2}^{2}.$ (11)
Here, $\alpha$ is the rate of transfer learning and $\mathbf{R}_{TL}$ is the (time-dependent) state of the reservoir (with the same number of time-snapshots as $\mathbf{X}_{TL}$). The update to the output layer and new output layer are given explicitly by
$\delta W$ $=\left(\mathbf{X}_{TL}\mathbf{R}_{TL}^{\prime}-W_{\mathrm{out}}\mathbf{R}_{TL}\mathbf{R}^{\prime}_{TL}\right)\left(\mathbf{R}_{TL}\mathbf{R}^{\prime}_{TL}+\alpha I\right)^{-1},$ (12)
@@
Transfer learning has been used with ESNs to successfully predict individual trajectories for shallow-water equations [11, 1] and the gKS equation [1]. In this paper, we show that transfer learning allows us to successfully track changes in the statistical signature of the gKS equation as either the size of the spatial domain or the dispersion relation vary.
@@ 4 Results
We focus on an important statistical feature of gKS dynamics, namely the power spectrum. The power spectrum is defined as the map from $\mathbb{Z}$ into $\mathbb{R}$ given by
@@
We first consider the KS equation and we vary the parameter $L$. Varying $L$ is equivalent to varying the number of linearly unstable Fourier modes. We show that for each of several values of $L$, the ESN accurately predicts the power spectrum of the KS equation with spatial domain $[0,L]$, once trained using KS trajectories from the same spatial domain. We then show that transfer learning is effective for various tuples $(L_{1},L_{2})$ of length parameters. That is, the ESN accurately predicts the power spectrum of the KS equation with parameter $L_{2}$ after a two-step training process. First, the ESN is trained using trajectories of the KS equation with parameter $L_{1}$. Second, the training is updated using a small amount of trajectory information from the KS equation with parameter $L_{2}$.
We conclude Section 4 by carrying out the program in the previous paragraph for the gKS equation. This time, we fix $L$ and vary the parameter $\gamma$ in the dispersion relation. We show that when $\gamma$ is small, the ESN accurately predicts power spectra and transfer learning is effective.
4.1 Resolution study
In [1, Section 5.1.3], Alam analyzed numerically how spatial resolution affects the performance of ESNs when using them to predict the power spectrum of the gKS equation. In particular, Alam considered $L=43$ and $\gamma=0.1$ and analyzed how numerically computed power spectra depend on the spatial resolution. In addition, Alam investigated how averaging time affects the accuracy of the computation of power spectra. Alam determined that spatial resolution $N_{x}=256$, averaging time $T=1000$ Lyapunov times, integration timestep $\delta t=0.001$, and snapshot sampling timestep $\Delta t=0.25$ are sufficient to accurately reproduce the power spectrum of the gKS equation in the chaotic regime. We therefore use these parameter values for the simulations in this paper, unless stated otherwise.
4.2 Predicting the power spectrum of the KS equation
4.2.1 Predicting the power spectrum of the KS equation without transfer learning
@@
$\lambda_{i}\left(L\right)\approx 0.093-0.94\left(i-0.39\right)/L.$ (15)
@@
We performed the following experiment for each value of $L$. We began with an untrained ESN. We generated $30$ KS trajectories of length $T=1000$ Lyapunov times using random initial data of the form
@@
Before considering prediction of the power spectrum, it is natural to ask how well the trained ESN predicts individual KS trajectories. Alam [1, Section 4.3.2] investigated this question. He found that ESN predictions of individual trajectories remain accurate up to approximately $7$ Lyapunov times. This timescale is consistent with results published earlier [45]. Alam [1, Section 4.3.2] found that the timescale for which predictions of individual trajectories remains accurate decays approximately linearly with $L$. As $L$ increases, one can potentially deal with this decay by training a larger ESN or by training a coupled network of ESNs. Nevertheless, we are unaware of rigorous mathematical guarantees associated with single-orbit prediction of spatiotemporally chaotic dynamics using ESNs. That is, we are unaware of rigorous lower bounds on the number of Lyapunov times for which the single-orbit prediction remains accurate. This motivates our focus on the power spectrum. The idea is that once the ESN can predict the power spectrum, it can do so for a very long time. In particular, it can continue to accurately predict the power spectrum long after single-orbit predictions have become inaccurate.
Figure 1 demonstrates that for each of the four values of $L$ in Table 1, the ESN accurately predicts the power spectrum of the KS equation (once trained). Each panel compares the power spectrum obtained by direct numerical simulation (DNS) to the ESN prediction. We average over $10$ trajectories used for prediction to obtain each DNS-based curve.
@@
Figure 1: Once trained on KS trajectories from spatial domain $[0,L]$, the ESN then accurately predicts the power spectrum of the KS equation with the same spatial domain. Each blue curve shows the power spectrum obtained from direct numerical simulation, averaged over $10$ trajectories used for prediction. Red curves show the ESN predictions of the power spectrum. Importantly, the ESN predictions of the power spectrum remain accurate long after ESN predictions of individual trajectories become inaccurate. Spatial domain sizes $L=22$ (top left), $L=29$ (top right), $L=35$ (bottom left), and $L=43$ (bottom right) correspond to four different numbers of linearly unstable Fourier modes (see Table 1). Plots show $\log(e_{k})$ versus $\log(k)$, where $k$ denotes wavenumber.
4.2.2 Predicting the power spectrum of the KS equation using transfer learning
When $L$ varies, it is computationally expensive to train a de novo ESN to predict the power spectrum for each new value of $L$. We therefore investigate the efficacy of transfer learning with respect to $L$ for the prediction of the power spectrum.
@@ 4.2.2 Predicting the power spectrum of the KS equation using transfer learning
We performed the following experiment. First, we trained the ESN in the $L=22$ regime using $20$ trajectories from that regime. Second, we set the transfer learning rate to $\alpha=0.005$ and then updated the training of the ESN using trajectory data from the $L=43$ regime. (We tested several values of $\alpha$, including $0.1$, $0.01$, $0.005$, $0.002$, and $0.001$. Empirically, $\alpha=0.005$ worked well.) We report the amount of TL data used for the training update as a percentage of the $20$ trajectories from the $L=43$ regime used for training.
In the absence of transfer learning, the ESN (trained only on trajectories from the $L=22$ regime) performs poorly in the $L=43$ regime for both single-orbit prediction and the prediction of the power spectrum [1]. However, transfer learning dramatically improves the prediction of the power spectrum in the $L=43$ regime. Figure 2 illustrates this dramatic improvement as a function of amount of TL data. The blue curve (same curve in each panel) is the power spectrum in the $L=43$ regime, computed by averaging over the $10$ DNS trajectories from this regime used for prediction. This curve serves as ground truth. The red curve (same curve in each panel) is the power spectrum for $L=22$ predicted by the ESN after having been trained on $20$ DNS trajectories from the $L=22$ regime. Comparing the red and blue curves, we see that the power spectrum for $L=22$ differs considerably from that for $L=43$. The green curves illustrate the ESN prediction of the power spectrum for the $L=43$ regime after transfer learning of level $10\%$ (left), $25\%$ (center), and $50\%$ (right). The quality of the prediction based on transfer learning rapidly improves as the level of TL data increases. In particular, the prediction is nearly perfect at level $50\%$ (right).

### v2

@@ Abstract
In this paper, we explore the predictive capabilities of echo state networks (ESNs) for the generalized Kuramoto-Sivashinsky (gKS) equation, an archetypal nonlinear PDE that exhibits spatiotemporal chaos. Our research focuses on predicting changes in long-term statistical patterns of the gKS model that result from varying the dispersion relation or the length of the spatial domain. We use transfer learning to adapt ESNs to different parameter settings and successfully capture changes in the underlying chaotic attractor. Previous work has shown that transfer learning can be used effectively with ESNs for single-orbit prediction. The novelty of our paper lies in our use of this pairing to predict the long-term statistical properties of spatiotemporally chaotic PDEs. Nevertheless, we also show that transfer learning nontrivially improves the length of time that predictions of individual gKS trajectories remain accurate.
Keywords: echo state networks, generalized Kuramoto-Sivashinsky equation, transfer learning
@@ 1 Introduction
In this paper, we explore the predictive capabilities of echo state networks (ESNs) in the context of spatiotemporal chaos. An echo state network is a simple reservoir computing architecture with the echo state property [30, 31, 42, 8, 32, 23, 65]. We utilize a methodology that integrates ESNs with transfer learning (TL) [29, 13].
It is often useful to model a physical process using a nonstationary dynamical system. Here, the dynamical model itself varies in time and may have time-dependent parameters. It is challenging to develop machine learning (ML) models that can track parametric changes in nonstationary dynamical systems. In particular, data in a new parameter regime might be missing, or insufficient to properly train a de novo ML model. There exists substantial recent work on developing ML models that can track such time-dependent parametric changes [48, 49, 52, 9, 38, 64].
In this paper, we integrate ESNs with transfer learning to predict changes in long-term statistical properties of spatiotemporally chaotic PDEs that result from instantaneously jumping from one parameter regime to another. Our work here complements existing work on developing ML models that can track parametric changes, as this existing work assumes parameters evolve continuously in time. The idea is that a well-trained ESN in the original parameter regime will maintain predictive power in the new parameter regime, provided the training of the ESN is updated with an additional amount of data from the new parameter regime. We now describe this idea mathematically.
@@
in the weak-$\ast$ topology, where $\delta_{z}$ denotes the Dirac-$\delta$ measure at $z$. In this setting, we ask two questions. First, if we fix a parameter value $\gamma_{1}\in\Gamma$ and train an echo state network using trajectories from $\Phi_{\gamma_{1}}$, can the echo state network then predict statistical properties of $(f\circ\Phi_{\gamma_{1}}^{t})_{t\geqslant 0}$, where $f:M\to\mathbb{R}^{n}$ is a given observable? Second, assuming an affirmative answer to the first question, how might we predict the statistical properties of $(f\circ\Phi_{\gamma_{2}}^{t})_{t\geqslant 0}$ with respect to a different parameter value $\gamma_{2}\in\Gamma$?
The point of this paper is to answer these two questions. Our approach is as follows. Suppose the echo state network has been trained using trajectories from $\Phi_{\gamma_{1}}$ and can successfully predict statistical properties of $(f\circ\Phi_{\gamma_{1}}^{t})_{t\geqslant 0}$. Since invariant measures of dynamical systems can depend sensitively on parameters, the invariant measure $\mu_{\gamma_{2}}$ associated with a different parameter value $\gamma_{2}$ might differ meaningfully from $\mu_{\gamma_{1}}$. Consequently, we would not expect the echo state network to successfully predict statistical properties of $(f\circ\Phi_{\gamma_{2}}^{t})_{t\geqslant 0}$ when trained using trajectories from $\Phi_{\gamma_{1}}$. We conjecture that the predictive capability of the echo state network can be ‘transferred’ from the $\gamma_{1}$-system to the $\gamma_{2}$-system using transfer learning [29, 13]. That is, the echo state network will successfully predict statistical properties of $(f\circ\Phi_{\gamma_{2}}^{t})_{t\geqslant 0}$ after we update its training using an additional new dataset from the new regime. The amount of new data depends on the dynamics of the underlying equations.
@@
As discussed above, we assume that the gKS model admits an ergodic invariant measure that describes the asymptotic distribution of typical initial data. We assume that the trained ESNs we use for prediction also possess this property. We have verified this hypothesis numerically.
Although this paper focuses on long-term statistical prediction, we nevertheless show that transfer learning nontrivially improves the length of time that predictions of individual gKS trajectories remain accurate. In particular, we report a gain of one to two Lyapunov times when using transfer learning to account for changes in the length of the spatial domain.
Transfer learning is a popular concept in many areas of machine learning, including speech recognition and image processing. However, transfer learning for echo state networks has received comparatively little attention in the literature. We conclude that transfer learning can allow echo state networks to maintain predictive power when model parameters instantaneously jump from one parameter set to another. Such flexibility is important when statistical properties of the dynamics depend meaningfully on model parameters.
@@ 2.1 Background
with the periodic boundary condition $u(x+L,t)=u(x,t)$ for all $x$ and every $t\geqslant 0$. The KS equation behaves nicely as an infinite-dimensional dynamical system. In particular, the KS model admits a finite-dimensional attractor and is therefore in some sense equivalent to a finite-dimensional dynamical system [15, 22, 46]. See [58] for an overview. Moreover, the KS model admits an inertial manifold that contains the global attractor [34, 14, 59]. The KS equation has been studied extensively, both analytically and computationally [37, 28, 56, 44].
The parameter $L$ (the length of the spatial domain) plays the role of a bifurcation parameter for the KS model. This parameter determines the number of linearly unstable Fourier modes, $S_{\mathrm{unst}}$. In addition, the maximum of the coefficient of the linearized equation occurs at $M_{\mathrm{unst}}$. We refer to $M_{\mathrm{unst}}$ as “the most unstable wavenumber”. These quantities are given by
@@
Here $\lfloor\cdot\rfloor$ denotes the greatest integer function.
@@ 3 Echo state networks and transfer learning
The input matrix $W_{\mathrm{in}}$ is generated component-wise from a uniform distribution on $\beta_{1},\beta_{1}]$. The adjacency matrix is generated as $A=\beta_{2}W_{0}/\lambda$, where $W_{0}$ is a sparse matrix (typically turning on less than $10\%$ of the possible connections), $\lambda$ denotes the largest eigenvalue of $W_{0}$, and $\beta_{2}$ is a scaling parameter.
@@
where $\mu$ is a regularization parameter, $\mathbf{X}$ is the training data (trajectories), and $\mathbf{R}$ denotes the reservoir timeseries (with $A$ and $W_{\mathrm{in}}$ fixed). $W_{\mathrm{out}}$ can be expressed explicitly as
@@
Here, $\mathbf{R}^{\prime}$ denotes the transpose of $\mathbf{R}$. Throughout this paper, we use ESN parameters $D=5000$ (same as in [50]), $\beta_{1}=0.01$, $\beta_{2}=0.1$, and $\mu=0.000005$, unless stated otherwise.
The ESN setup we use in this paper is similar to that of [13]. The common approach is that the values of $\beta_{1}$ and $\beta_{2}$ should be chosen so that $Ar(t)+W_{\mathrm{in}}X(t)$ evolves in the nonlinear regime of $\tanh(\cdot)$ and the ESN operates in the unsaturated regime. Choosing appropriately depends upon the problem at hand. We performed a basic search to determine optimal hyperparameter values. Our search revealed that the ESN performs considerably suboptimally when $\beta_{2}\lesssim 1$ and $\beta_{1}\in[0.1,0.5]$. The performance of the ESN is not very sensitive to hyperparameter values as long as $\beta_{1}\in[0.01,0.1]$ and $\beta_{2}\in[0.1,0.4]$. The value of $\mu$ has been taken from [12]. Since our study involves long stationary simulations where initial transient behavior is not important, we do not use reservoir warm-up in this work. We performed a synchronization test (not included here for the sake of brevity), showing that the reservoir synchronizes with 2 to 4 timesteps.
Transfer learning [29, 13] is a machine-learning approach wherein knowledge in a source domain is used to improve predictive performance in a target domain. In our context, the source and target domains are simply two sets of gKS parameters. Our central premise for linking transfer learning to ESNs is the following. Once the ESN is trained well in the source domain, it will retain predictive power in the target domain after receiving a training update using a relatively small transfer-learning dataset from the target domain. However, the amount of transfer learning data might depend on a particular dynamic regime.
Writing $\mathbf{X}_{\mathrm{TL}}$ for the transfer-learning dataset, the update to the output layer is obtained by solving the optimization problem given by
$\delta W=\operatorname*{arg\,min}_{\delta W}\left\|(W_{\mathrm{out}}+\delta W)\mathbf{R}_{\mathrm{TL}}-\mathbf{X}_{\mathrm{TL}}\right\|_{2}^{2}+\alpha\|\delta W\|_{2}^{2}.$ (11)
Here, $\alpha$ is the rate of transfer learning and $\mathbf{R}_{\mathrm{TL}}$ is the (time-dependent) state of the reservoir (with the same number of time-snapshots as $\mathbf{X}_{\mathrm{TL}}$). The update to the output layer and new output layer are given explicitly by
$\delta W$ $=\left(\mathbf{X}_{\mathrm{TL}}\mathbf{R}_{\mathrm{TL}}^{\prime}-W_{\mathrm{out}}\mathbf{R}_{\mathrm{TL}}\mathbf{R}^{\prime}_{\mathrm{TL}}\right)\left(\mathbf{R}_{\mathrm{TL}}\mathbf{R}^{\prime}_{\mathrm{TL}}+\alpha I\right)^{-1},$ (12)
@@
In this paper, we use $\alpha=0.005$. We performed a grid search to determine the optimal value of this hyperparameter. In particular, we tested several values of $\alpha$, including $0.1$, $0.01$, $0.005$, $0.002$, and $0.001$. Empirically, $\alpha=0.005$ works well.
Transfer learning has been used with ESNs to successfully predict individual trajectories for shallow-water equations [13, 1] and the gKS equation [1]. In this paper, we show that transfer learning allows us to successfully track changes in the statistical signature of the gKS equation when either the size of the spatial domain or the dispersion relation instantaneously jumps from one regime to another.
@@ 4 Results
We focus on an important statistical feature of gKS dynamics, namely the power spectrum. The time-averaged power spectrum is defined as the map from $\mathbb{Z}$ into $\mathbb{R}$ given by
@@
We first consider the KS equation and we vary the parameter $L$. Varying $L$ is equivalent to varying the number of linearly unstable Fourier modes. For each of several values of $L$, we train the ESN using trajectories of the KS equation with spatial domain $[0,L]$. We show that the trained ESN accurately predicts the power spectrum of the KS equation when the spatial domains used for training and prediction are the same.
We then show that transfer learning is effective for various tuples $(L_{1},L_{2})$ of length parameters. That is, the ESN accurately predicts the power spectrum of the KS equation with parameter $L_{2}$ and the correlation dimension of the corresponding attractor after a two-step training process. First, the ESN is trained using trajectories of the KS equation with parameter $L_{1}$. Second, the training is updated using some trajectory information from the KS equation with parameter $L_{2}$.
We complete our suite of experiments for the KS equation by showing that transfer learning (with respect to $L$) nontrivially improves the length of time that predictions of individual KS trajectories remain accurate. In particular, we report a gain of one to two Lyapunov times using a modest percentage of data for transfer learning ($5\%$).
We conclude Section 4 by carrying out the program described above for the gKS equation. This time, we fix $L$ and vary the parameter $\gamma$ in the dispersion relation. We show that when $\gamma$ is small, the well-trained ESN accurately predicts the power spectrum and the correlation dimension of the attractor, when the values of $\gamma$ used for training and prediction match. Further, we show that transfer learning is effective in the small-$\gamma$ regime.
Parameter selection for experiments involving the power spectrum. In the previous study [1, Section 5.1.3], we determined that $N_{x}=256$, averaging time $T=10000$ (approximately $1000$ Lyapunov times), integration timestep $\delta t=0.001$, and snapshot sampling timestep $\Delta t=0.25$ are sufficient for resolving power spectra. We use these parameters for the experiments in this paper, unless specifically stated otherwise. Since we always use $N_{x}=256$ points in for the discretization, $\Delta x$ depends on $L$.
4.1 Predicting the statistical behavior of the KS equation without transfer learning
@@
$\lambda_{i}(L)\approx 0.093-\frac{(0.94)(i-0.39)}{L}.$ (15)
@@
We performed the following experiment for each value of $L$. We began with an untrained ESN. We generated $30$ KS trajectories of length $T=10000$ (approximately $1000$ Lyapunov times) using random initial data of the form
@@
Figure 1 demonstrates that for each of the four values of $L$ in Table 1, the ESN accurately predicts the power spectrum of the KS equation (once trained). Each panel compares the power spectrum obtained by direct numerical simulation (DNS) to the ESN prediction. We average over $10$ DNS trajectories used for prediction to obtain each DNS-based curve.
@@
Figure 1: Once trained on KS trajectories from the spatial domain $[0,L]$, the ESN then accurately predicts the time-averaged power spectrum of the KS equation with the same spatial domain. Each blue curve shows the power spectrum obtained from direct numerical simulation, averaged over $10$ DNS trajectories used for prediction. Red curves show the ESN predictions of the power spectrum. Importantly, the ESN predictions of the power spectrum remain accurate long after ESN predictions of individual trajectories become inaccurate. Spatial domain sizes $L=22$ (top left), $L=29$ (top right), $L=35$ (bottom left), and $L=43$ (bottom right) correspond to four different numbers of linearly unstable Fourier modes (see Table 1). Plots show $\log(e_{k})$ versus $\log(k)$, where $k$ denotes wavenumber.
4.2 Predicting the statistical behavior of the KS equation when the length of the spatial domain instantaneously increases using transfer learning
When $L$ changes instantaneously, it is computationally expensive to train a de novo ESN to predict the power spectrum for each new value of $L$ because one needs to generate a new training dataset for each new such value. Thus, in the context of using ESNs for systems with different values of parameters, generating new training data can be a computational bottleneck. Transfer learning can potentially reduce the computational cost by utilizing smaller datasets for retraining. In this section, we demonstrate the efficacy of transfer learning with respect to $L$ for the prediction of the power spectrum. Further, we show that when $L$ increases instantaneously, transfer learning restores the ability of the ESN to predict the correlation dimension of the attractor.
@@
We performed the following experiment. First, we trained the ESN in the $L=22$ regime using $20$ trajectories from that regime. Second, we set the transfer learning rate to $\alpha=0.005$ and then updated the training of the ESN using trajectory data from the $L=43$ regime. We report the amount of TL data used for the training update relative to the size of the dataset used to train the ESN when $L=43$ in the experiment from Section 4.1 ($20$ trajectories).